A pool with dimensions  5 m long,4 .5m wide and 2,m deep has water to a depth of 1/3m . How many cubic  meter of water must be added to make pool full

Grade:

1 Answers

Meenakshy Satheesh
9 years ago
The pool is cuboidal obviously...
Length of the pool, l = 5m
Breadth of the pool, b = 4.5m
Height of the pool, h = 2m
The total water needed to completely fill the pool = Volume of the pool
                                                                       = l*b*h
                                                                       =5*4.5*2
                                                                       =45 m3
Given, that 1/3rd of the pool is filled, i.e., 45*1/3 = 15m3
So, amount of water to be filled = 45m3 – 15m3
                                              = 30m3.